STRUCTURAL STUDIES OF CH_{3}SiF_{2}-X (X = NCO, Cl) BY MICROWAVE SPECTROSCOPY

GAMIL A. GUIRGIS, KORREDA K. GAUSE, Department of Chemistry & Biochemistry, College of Charleston, Charleston, SC 29424 USA; \underline{NATHAN A. SEIFERT}, DANIEL P. ZALESKI, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904-4319; MICHAEL H. PALMER, School of Chemistry, University of Edinburgh, West Mains Road, Edinburgh EH9 3JJ, UK; REBECCA A. PEEBLES, SEAN A. PEEBLES, LENA F. ELMUTI, DANIEL A. OBENCHAIN, Department of Chemistry, Eastern Illinois University, 600 Lincoln Avenue, Charleston, IL, 61920 USA.

The structures of CH_{3}SiF_{2}-NCO and CH_{3}SiF_{2}-Cl have been studied by molecular rotational spectroscopy in the 6.5-18 GHz band. The rotational spectrum was measured by cavity Fourier transform microwave (FTMW) and chirpedpulse FTMW spectroscopy. The experiment targeted the study of CH_{3}SiF_{2}-NCO, but CH_{3}SiF_{2}-Cl was also observed as an impurity. Due to the dynamic range achieved on these spectra, all isotopologs with natural abundance ≥0.2% were assigned, which includes two doubly-substituted isotopologs for the chloride (^{29}Si/^{37}Cl and ^{30}Si/^{37}Cl). Strategies for obtaining the molecular structure for these two molecules using either a Kraitchman analysis (to obtain a partial substitution structure) or r_{0} analysis (with additional constraints on the structure supplied by the theoretical structure) will be discussed. Derived structural parameters for the CH_{3}-SiF_{2}-X base structure are the same for the two compounds. The hyperfine and internal rotation effects in the spectra have been analyzed for all isotopologs and the Hamiltonian parameters are in very good agreement with ab initio results. The barriers to methyl group internal rotation for the two compounds 446(50) cm^{−1} and 463(3) cm^{−1} and are independent of the isotopic structure of the heavy atom frame.